A solution is made using 200.0 mL of methanol (density 0.792 g/mL) and 301.1 mL of water (density 1.000 g/mL). What is the mass percent of the methanol?
Answer #1

density \(=\frac{\text { mass }}{\text { volume }} \)

\(0.792=\frac{m}{200.0}\)

mass of methanol \(=200.0 \times 0.792\)

\(=158.49\)

mass of water \(=301.1 \times 1.0\)

\(=301.19\)

Total mass \(=158.4+30101\)

Mass % of methanol \(=\frac{\text {( mass of methanol) }}{\text { total mass }} \times 100\)

\(=\frac{158.4}{459.5} \times 100\)

\(=34.47 \%\)

answered by: BEN